# TriDF: Triplane-Accelerated Density Fields for Few-Shot Remote Sensing Novel View Synthesis

### v1

@@ IV-D1 Ablation of Model Structure
We conducted comprehensive ablation experiments on our proposed hybrid representation to validate the critical design elements for efficient and accurate novel view synthesis. Our evaluation focused on three key components: the triplane representation, density fields, and the image-based rendering framework. To assess the hybrid representation thoroughly, we developed two alternative implementations for comparative analysis, as illustrated in Fig. 5. These variations include the aggregation of reference features in the triplane branch and the separation of feature interactions between triplane and density fields branches.
The pre-trained encoder demonstrates capability in capturing prior information about scene appearance and geometry. Intuitively, incorporating such priors into our scene representation (whether in the triplane or density fields) could enhance the model’s spatial understanding. However, our experiments in Table III reveal unexpected limitations when implementing the IBR framework in the triplane representation, as shown in Fig. 5(a). Specifically, the triplane representation fails to learn detailed scene information due to the nature of projection transformations: the triplane features of any spatial point maintain a one-to-one correspondence with the reference image features. Given that the encoder remains frozen, the reference image features become constant. Consequently, the triplane features tend to overfit the reference features and fail to update. Based on these findings, we try to integrate IBR features into the density fields branch. Experimental results demonstrate that IBR features can provide beneficial prior information about scene structures for the density fields. In comparison to our other implementations, our method is capable of reconstructing more intricate and precise geometry, as illustrated in Fig. 6.
@@
As illustrated in Fig. 5(b), we explore modeling color and volume density as entirely independent components. While this implementation successfully generated high-quality images (seeing Fig. 6), our final method introduces an interaction between the triplane representation and density fields. Specifically, the density fields generate supplementary features alongside volume density, which are then concatenated with triplane features for color prediction. This interactive approach demonstrates superior view synthesis quality compared to the independent modeling approach, without introducing additional computational complexity to the density fields.
@@ VI Conclusion
In this work, we explore the potential of triplane representation in remote sensing novel view synthesis and propose an efficient hybrid 3D representation, Triplane-Accelerated Density Fields (TriDF). Our approach achieves superior rendering quality from sparse viewpoint inputs while achieving a faster reconstruction speed (30$\times$) compared to NeRF-based methods. We improve the training efficiency by decoupling color and volume density, leveraging the triplane representation to model high-frequency color information while utilizing the density fields to capture continuous volume density. Furthermore, we introduce an efficient depth-guided optimization technique to prevent overfitting issues in few-shot NVS. Experiments conducted across multiple scenes demonstrate that our approach outperforms other state-of-the-art few-shot NVS methods while maintaining fast reconstruction within only 5 minutes. Consequently, TriDF exhibits immense potential for 3D interpretation in remote sensing applications.

### v2

@@ IV-D1 Ablation of Model Structure
We conducted comprehensive ablation experiments on our proposed hybrid representation to validate the critical design elements for efficient and accurate novel view synthesis. Our evaluation focused on three key components: the triplane representation, density fields, and the image-based rendering framework. To assess the hybrid representation thoroughly, we developed two alternative implementations for comparative analysis, as illustrated in Fig. 5. These variations include the aggregation of reference features in the triplane branch and the separation of feature interactions between the triplane and density fields branches.
@@
The pre-trained encoder demonstrates the capability of capturing prior information about scene appearance and geometry. Intuitively, incorporating such priors into our scene representation (whether in the triplane or density fields) could enhance the model’s spatial understanding. However, our experiments in Table III reveal unexpected limitations when implementing the IBR framework in the triplane representation, as shown in Fig. 5(a). Specifically, the triplane representation fails to learn detailed scene information due to the nature of projection transformations: the triplane features of any spatial point maintain a one-to-one correspondence with the reference image features. Given that the encoder remains frozen, the reference image features become constant. Consequently, the triplane features tend to overfit the reference features and fail to update. Based on these findings, we try to integrate IBR features into the density fields branch. Experimental results demonstrate that IBR features can provide beneficial prior information about scene structures for the density fields. In comparison to our other implementations, our method is capable of reconstructing more intricate and precise geometry, as illustrated in Fig. 6.
@@
As illustrated in Fig. 5(b), we explore modeling color and volume density as entirely independent components. While this implementation successfully generated high-quality images, as shown in Fig. 6, our final method introduces an interaction between the triplane representation and density fields. Specifically, the density fields generate supplementary features alongside volume density, which are then concatenated with triplane features for color prediction. This interactive approach demonstrates superior view synthesis quality compared to the independent modeling approach, without introducing additional computational complexity to the density fields.
@@ IV-D2 Ablation of Optimization Strategy
Figure 9: Comparative visualization of different scene types. We report the average PSNR on test views to evaluate the model’s performance in different scenes.
@@ IV-D3 Impacts of Model Parameters
IV-D4 Performance on Different Scenes
We also evaluate the model’s performance across different types of scenes. As illustrated in the Fig. 9, we roughly categorize the influencing factors into three parts: ground object types, the number of buildings, and building height. First, scenes with high natural vegetation coverage (e.g., Mountain and Park) exhibit better reconstruction quality than scenes with dense building coverage. This is because natural vegetation tends to have simpler and more tolerant texture structures, making them less sensitive to errors. In contrast, buildings contain fine-grained geometric details, where even small deviations can lead to a noticeable degradation in synthesis quality. The number of buildings directly reflects the complexity of a scene. Scenes with fewer buildings (e.g., Building #1 and College) provide more complete view coverage, resulting in higher-quality synthesis. Conversely, scenes with many buildings (such as Church) are more difficult to reconstruct due to their intricate structures and the occlusions between buildings. Finally, due to the limited number of viewpoints and restricted shooting angles, taller buildings introduce more severe occlusion. As a result, scenes with relatively flat terrain (e.g., Factory and School) tend to achieve better reconstruction quality.
@@ VI Conclusion
In this work, we explore the potential of triplane representation in remote sensing novel view synthesis and propose an efficient hybrid 3D representation, Triplane-Accelerated Density Fields (TriDF). Our approach achieves superior rendering quality from sparse viewpoint inputs while achieving a faster reconstruction speed compared to NeRF-based methods. We improve the training efficiency by decoupling color and volume density, leveraging the triplane representation to model high-frequency color information while utilizing the density fields to capture continuous volume density. Furthermore, we introduce an efficient depth-guided optimization technique to prevent overfitting issues in few-shot NVS. Experiments conducted across multiple scenes demonstrate that our approach outperforms other state-of-the-art few-shot NVS methods while maintaining fast reconstruction within only 5 minutes. Consequently, TriDF exhibits immense potential for 3D interpretation in remote sensing applications.